OpenStudy (anonymous):

The area of a rectangle is represented by the polynomial function, A(x) = 2x^2 + 6x + 4. Find polynomial expressions that might represent the dimensions (length and width) of the rectangle.

OpenStudy (anonymous):

Perimeter of rectangle= sum of all 4 sides or 2(L+W) Width: 'x' Length: 2x+2 Perimeter= 40 2(x+2x+2)=40 "OR" x+x+2x+2+2x+2=40 6x+4=40 Subtract 4 from both sides to get x term alone on one side 6x=36 Divide both sides by 6 to get x alone x=6 Length= 14m

OpenStudy (radar):

The area is the product of the length and width. Do you agree?

OpenStudy (radar):

If you do, then by factoring the Area function A(x) getting to polynomials that would represent the length and width.

OpenStudy (radar):

What do you have for answer options?

OpenStudy (anonymous):

I have no clue what the answer is and I'm super confused. My brother told me the answer was (2x+2)(x+2)

OpenStudy (anonymous):

you must factor the polynomial in order to figure the expressions of each dimension. Since your question does not ask for any specific value,the answer can be expressed as